# Dynamical scaling study of three-dimensional XY spin glass toward the spin-chirality decoupling picture

### v1

@@ IV Estimation of Dynamical Exponent to Analyze the Critical Phenomena in SG Systems
Using the transition temperatures obtained from the analysis in Sec. III, we can compute the correlation length in a manner similar to that described in Sec. IV.1 and IV.2. The results of the correlation function computed at the transition temperatures are shown in Fig. 11. We subsequently perform a dynamical scaling analysis based on Eq. (16). In an optimization procedure, we discarded short-range correlations ($r<3$ for $f_{\mathrm{SG}}$ and $r<4$ for $f_{\mathrm{CG}}$) and data of small values ($f_{\mathrm{SG}}(r,t)<10^{-5}$ and $f_{\mathrm{CG}}(r,t)<2\times 10^{-6}$). The initial quantity of $\{\xi(t)\}$ was set to the values of correlations obtained by the second moment method, represented by Eq. (17) for SG and Eq. (18) for CG. The scaling plots by inference correlation length are shown in Fig. 12. Figure 13 shows the scaling result based on the dynamical scaling law in Eq. (16).
@@
From the inference results and Eq. (1), the value of $z$ is evaluated as $z=5.86(1)$ for the SG case and $z=5.24(7)$ for the CG case. The errors were assessed through a bootstrap[16] method with 10,000 trials. These results and the previously mentioned dynamical scaling analysis in Sec. III can be used to estimate the critical exponents. These results are compared with those of previous studies in Tables 1 and 2.
@@
The obtained results show improved accuracy for both the transition temperature and the critical exponents compared with previously reported findings. The dynamical exponent $z$ was found to be lower in the CG case than in the SG case, which supports the result that finite-size effects are more pronounced in the CG case (Fig. 1).
@@
Figure 11: The dynamical SG and CG correlation functions of the 3D ±$\pm J$ XY model, at the critical temperature for various Monte Carlo steps $t$: (a) analysis of the SG transition and (b) analysis of the CG transition.
@@
Figure 12: The scaling plots of the SG and CG correlation functions in the 3D ±$\pm J$ XY model: (a) analysis of the SG transition and (b) analysis of the CG transition.
Figure 13: The SG and CG correlation lengths in the 3D ±$\pm J$ XY model.
@@ V Summary and Discussion
We primarily aimed to validate a method for calculating the dynamical exponent from the correlation function, thereby enabling an analysis of SG systems. The obtained dynamical exponent appeared to be reliable, showed good accuracy, and was not limited to SG systems. The results confirm the reliability of this method. Using this approach, we analyzed the 3D ±$\pm J$ XY model. Compared with conventional methods, our method achieved a higher precision in evaluating the dynamical exponent. In addition, our results support the validity of the spin chirality decoupling picture in SG systems.
@@
The correlations estimated using our method also face problems related to finite-time corrections. Because we do not extrapolate time values for the calculation of the dynamical exponent, our proposed method is difficult to apply to regions with strong fluctuations. An example of this situation is the analysis of SG transitions near multi-critical points in the 3D ±$\pm J$ Ising model[2]. Making the method more widely applicable will require the development of an extrapolation method for $t\rightarrow\infty$. A reliable and systematic extrapolation scheme has recently been proposed in the NER method[19]. We expect that using such a scheme will lead to solutions to such problems. This work is also a future task.

### v3

@@ IV.2 Estimation of critical exponents
The results of the correlation function computed at the transition temperatures are shown in Fig. 12. We subsequently perform a dynamical scaling analysis based on Eq. (18). In an optimization procedure, we discarded short-range correlations ($r<3$ for $f_{\mathrm{SG}}$ and $r<4$ for $f_{\mathrm{CG}}$) and data of small values ($f_{\mathrm{SG}}(r,t)<10^{-5}$ and $f_{\mathrm{CG}}(r,t)<2\times 10^{-6}$). The initial values of $\{\xi(t)\}$ were set to those of correlations obtained by the second moment method, represented by Eq. (19) for SG and Eq. (20) for CG. The scaling plots by inference correlation length are shown in Fig. 13. Figure 14 shows the scaling result based on the dynamical scaling law in Eq. (18).
@@
From the inference results and Eq. (1), the value of $z$ is evaluated as $z=5.86(1)$ for the SG case and $z=5.24(7)$ for the CG case. The uncertainties were quantified as the standard deviations of the bootstrap distributions derived from 10,000 resamplings[14]. These results and the previously mentioned dynamical scaling analysis in Sec. IV.1 can be used to estimate the critical exponents. These results are compared with those of previous studies in Tables 1 and 2.
@@
The obtained results show improved accuracy for both the transition temperature and the critical exponents compared with previously reported findings. Here, the value of $\gamma/z\nu$ is found to be almost the same as those reported in previous studies employing NER analysis for the SG transition, whereas this is not the case for the CG transition[4, 9]. This discrepancy is likely due to the fact that those earlier studies did not take into account the temperature dependence of the dynamical exponent $z$ in the CG analysis. The dynamical exponent $z$ was found to be lower in the CG case than in the SG case, which supports the result that finite-size effects are more pronounced in the CG case (Fig. 7).
@@
Figure 12: The dynamical SG and CG correlation functions of the 3D ±$\pm J$ XY model at the critical temperature for various Monte Carlo steps $t$: (a) analysis of the SG transition and (b) analysis of the CG transition.
@@
Figure 13: The scaling plots of the SG and CG correlation functions in the 3D ±$\pm J$ XY model: (a) analysis of the SG transition with $\eta_{\mathrm{eff}}=0.6865$, and (b) analysis of the CG transition with $\eta_{\mathrm{eff}}=1.169$.
Figure 14: The SG and CG correlation lengths in the 3D ±$\pm J$ XY model. The dotted line indicates the slope corresponding to the estimated dynamic exponent $z$.
@@ V Summary and Discussion
We primarily aimed to validate a method for calculating the dynamical exponent from the correlation function, thereby enabling an analysis of SG systems. The obtained dynamical exponent appeared to be reliable, showed good accuracy, and was not limited to SG systems. The results confirm the reliability of this method. Using this approach, we analyzed the 3D ±$\pm J$ XY model. Compared with conventional methods, our method achieved a higher precision in evaluating the dynamical exponent. Through this analysis, we also examined the temperature dependence of $z(T)$ in the CG transition. In addition, our results support the validity of the spin chirality decoupling picture in SG systems.
@@
The correlations estimated using our method also face problems related to finite-time corrections. Because we do not extrapolate the scaling regime as $t\rightarrow\infty$ for the calculation of the dynamical exponent, our proposed method would suffer a difficult to apply to cases with strong fluctuations. An example of such a situation is the analysis of SG transitions near multi-critical points in the 3D ±$\pm J$ Ising model[2]. Making the method more widely applicable will require the development of an extrapolation method for $t\rightarrow\infty$. A reliable and systematic extrapolation scheme has recently been proposed in the NER method[19]. We expect that using such a scheme will lead to solutions to such problems. This work is also a future task.